OpenStudy (anonymous):

ok so I have NO clue how to do these!!! 7/8x-1=15/16x+4

OpenStudy (anonymous):

just bring the 7/8x over to the other side, and bring the 4 over to the side that the 7/8x was on. you get -5 = 1/16x or x = -80

OpenStudy (anonymous):

well that didnt work for the answer.